Quinine, C20H24N2O2 is the bitter tasting compound in tonic water. It is also the compound used to prevent malaria.

What is the mass of 0.25 mol quinine?


1
Expert's answer
2021-02-04T07:05:10-0500
"m=n*M"

M (C20H24N2O2) = 324 g/mol


"m=0.25*324=81 g"